# Covariances for Free: Exploiting Mean Distributions for Federated Learning with Pre-Trained Models

### v1

@@ Appendix C Proof of Proposition 2
By applying the variance definition and proposition 1, we obtain:
@@
Now, by considering the right term in Eq. 15, since $\hat{\mu}$ is an unbiased estimator of the population mean, then $\mathbb{E}[\hat{\mu}]=\mu$. Moreover, since we assume that the features for a single class across clients are i.i.d at initialization, we can re-use Proposition 1 by considering the all class features as a random sample of size $N$ from a population with mean $\mu$ and variance $\Sigma$. Consequently, the global sample mean $\hat{\mu}$ is has variance $\mathrm{Var}[\hat{\mu}]=\frac{\Sigma}{N}$. Then
@@
By replacing equations 16 and 17 in Eq. 15, and recalling that $N=\sum_{k=1}^{K}n_{k}$, we obtain:
@@ Appendix D Proof of Proposition 3
The first part, regarding Eq. 18, follows directly. From the ridge regression solution, $B=FY$, which is obtained by summing the features for each class and arranging them into the columns of a matrix. This results in the product of class means and samples per class.
@@
Replacing the global covariance $\hat{S}$ in Eq. 20, we obtain the final expression for $G$ as:
@@
Appendix E Communication Costs
@@
When computing communication costs we consider that the pre-trained models are on the clients and do not need to be communicated. We do not include cost of backward communication of classifier parameters from server to clients, since it is the same for all methods but is necessary if the models are fine-tuned after classifier initialization. All parameters are considered to be 32-bit floating point numbers (i.e. 4 bytes).
Appendix F Implementation Details
@@
Here, we provide details on learning rate (lr) used for all fine-tuning experiments with FedAdam. For ImageNet-R, we use a lr of 0.0001 for both server and clients for FedNCM, Fed3R and FedCOF initializations. For CUB200 and Cars, we use a server lr of 0.00001 and client lr of 0.00005 for Fed3R and FedCOF, while for FedNCM, we use a higher lr of 0.0001 for clients.
@@
For the linear probing experiments with FedAvg, we use a client lr of 0.01 and server lr of 1.0 for FedNCM across all datasets. For Fed3R and FedCOF initializations, we use a client lr of 0.001 and a server lr of 1.0 for all datasets. We use an Nvidia RTX 6000 GPU for our experiments.
Appendix G The FedCOF Oracle (Sharing Full Covariances)
@@
Similar to (Luo et al., 2021), we aggregate the class covariances from clients as follows:
@@
We use the aggregated class covariance from Eq. 21 and apply shrinkage to obtain $\hat{\Sigma}_{c}+\gamma I_{d}$ and use it in Eq. 11 for the oracle setting of FedCOF.
Appendix H Additional Experiments
Sampling multiple class means. We perform multiple class means sampling per client using ImageNet-R and show in Fig. 13(a) that using FedCOF with more class means shared from each client improves the performance. We also show in Fig. 13(b) the total number of means used per class on an average in Fig. 13(a) to perform the covariance estimation. The number of means used is less than the total number of clients due to the class-imbalanced or dirichlet distribution used to sample data for clients.
Figure 13: (a) Analysis of FedCOF performance with multiple class means per client on ImageNet-R. (b) Total number of means per class on average that are used to estimate the covariance for FedCOF in Fig. 13(a).
Linear probing after initialization experiments. We show in Fig. 15 that linear probing after FedCOF classifier initialization improves the accuracy significantly compared to FedNCM and is marginally better than Fed3R initialization across three datasets using SqueezeNet.
Figure 15: Analysis of the performance with linear probing using FedAvg (McMahan et al., 2017).
Comparison of training-free methods with linear probing. We also compare with our approach with the training-based federated linear probing without any initialization (where we perform FedAvg and learn only the classifier weights of models) and show in table Table 3 that FedCOF is more robust and communication-efficient compared to federated linear probing across several datasets. We follow the same settings as in Table 2. For first 4 datasets, we perform federated linear probing for 200 rounds with 30 clients per round using FedAvg with a client learning rate of 0.01. For iNat-120k, we train more for 5000 rounds.
Table 3: Comparison of different training-free methods using SqueezeNet with training-based Fed-LP (federated linear probing with FedAvg (McMahan et al., 2017) starting with pre-trained model and random classifier initialization) across 5 random seeds. FedNCM, Fed3R and the proposed FedCOF does not involve any training. We show the total communication cost (in MB) from all clients to server. The best results from each section are highlighted in bold.
CIFAR100 ImageNet-R CUB200 CARS iNat-120K
@@ Appendix H Additional Experiments
Fed-LP 59.9$\pm$0.2 2458 37.8$\pm$0.3 4916 46.8$\pm$0.8 4916 33.1$\pm$0.1 4817 28.0$\pm$0.6 1.6×$\times 10^{6}$
FedNCM 41.5$\pm$0.1 5.9 23.8$\pm$0.1 7.1 37.8$\pm$0.3 4.8 19.8$\pm$0.2 5.4 21.2$\pm$0.1 111.8
Fed3R 56.9$\pm$0.1 110.2 37.6$\pm$0.2 111.9 50.4$\pm$0.3 109.6 39.9$\pm$0.2 110.2 32.1$\pm$0.1 9837.3
FedCOF (Ours) 56.1$\pm$0.2 5.9 37.8$\pm$0.4 7.1 53.7$\pm$0.3 4.8 44.0$\pm$0.3 5.4 32.5$\pm$0.1 111.8
Experiments with ResNet18. We perform experiments with pre-trained ResNet18 in table Table 4. For FedAvg and FedAdam, we train for 200 rounds with 30 clients per round. For FedAvg, we train with a client learning rate of 0.001 and server learning rate of 1.0. For FedAdam, we train with a client learning rate of 0.001 and a server learning rate of 0.0001.
Table 4: Comparison of different training-free methods using pre-trained ResNet18 for 100 clients with training-based federated learning baselines FedAvg (McMahan et al., 2017) and FedAdam (Reddi et al., 2020) starting from a pre-trained model. We train for 200 rounds for FedAvg and FedAdam. FedNCM, Fed3R and the proposed FedCOF do not involve any training. We also show the performance of fine-tuning with FedAdam after classifier initialization. For fine-tuning experiments we only train for 100 rounds after initialization. We show the total communication cost (in MB) from all clients to server. The best results from each section are highligted in bold.
CIFAR100 ImageNet-R

### v3

@@ Appendix C Proof of Proposition 1
By applying the variance definition, along with the expectation and variance of the sample mean (see Equation˜10), we obtain:
@@
Now, by considering the right term in Equation˜13, since $\hat{\mu}$ is an unbiased estimator of the population mean, then $\mathbb{E}[\hat{\mu}]=\mu$. Moreover, since we assume that the features for a single class across clients are i.i.d at initialization, we can use re-use the result in Equation˜10 by considering the all class features as a random sample of size $N$ from a population with mean $\mu$ and variance $\Sigma$. Consequently, the global sample mean $\hat{\mu}$ is has variance $\mathrm{Var}[\hat{\mu}]=\frac{\Sigma}{N}$. Then
@@
By using Equation˜14 and Equation˜15 in Equation˜13, and recalling that $N=\sum_{k=1}^{K}n_{k}$, we obtain:
@@ Appendix D Proof of Proposition 2
The first part, regarding Equation˜16, follows directly. From the ridge regression solution, $B=FY$, which is obtained by summing the features for each class and arranging them into the columns of a matrix. This results in the product of class means and samples per class.
@@
Replacing the global covariance $\hat{S}$ in Equation˜18, we obtain the final expression for $G$ as:
@@
Appendix E Sampling Multiple Class Means per Client
From a theoretical perspective, although the estimator $\hat{\Sigma}_{c}$ is unbiased in Equation˜5, this only ensures that its expected value equals the true covariance $\Sigma_{c}$ – not that any individual estimate is accurate. Its variance still depends on the number of independent means $K$ sampled – that is the number of clients in the federation $K$ which contains class $c$.
Therefore, increasing $K$ leads to a tighter concentration of $\hat{\Sigma}_{c}$ around its expectation, reducing the overall mean square error (MSE) – the sum of variance and squared bias (which remains zero since the estimator is unbiased) – between the estimator and the true covariance matrix. Moreover, where the number of clients is small relative to the feature dimension $d$, the estimate may be ill-conditioned. The shrinkage term $\gamma I_{d}$ in Equation˜6 improves numerical stability in these settings, trading a small amount of bias for a significant reduction in variance – even when only a few clients are available.
Sampling strategy. We propose using a simple multiple mean sampling strategy following sampling without replacement. We consider the number of means $\mathcal{M}$ to sample as a fixed number which is a hyperparameter. For each class, we take disjoint random sets from $n_{k,c}$ samples in a client and compute the mean for these subsets. We take disjoint sets to avoid computing similar sample means. The only condition we enforce is that clients use atleast 2 samples to compute a mean. If a client does not have atleast $2\mathcal{M}$ samples, we send less than $\mathcal{M}$ sample means. For instance, if a client has 3 samples, we compute and share a single mean.
We observe that the proposed sampling approach improves performance. However, more sophisticated sampling approaches could be employed if the user is interested in improving performance in FL settings with very few clients. One approach could be sampling means based on the number of samples $n_{k,c}$ instead of using a fixed number of means to sample from every client. If a client has more samples, it could share more sample means and this would thus share more means overall and improve the covariance estimates. Future work could optimize this multiple mean sampling approach to better suit fewer client FL settings.
Appendix F On Excluding Between-class Scatter
Intuitively, to represent the feature distribution of each class we do not really need to consider the relationships between different classes which represent the distribution of the overall dataset since our goal is to estimate the class-specific classifier weights using these covariances. Based on this intuition, we propose to remove the between-class scatter from Equation˜7 and initialize the classifier weights using only the respective within-class scatter matrices.
We also analyze in Table˜6 using the centralized setting how the different scatter matrices affect overfitting of the model. We observe that, while both methods achieve similarly high training accuracies, Ridge Regression consistently underperforms on the test set. This suggests that incorporating $G_{\text{btw}}$ introduces an overfitting effect, as the classifier learns directions that do not transfer well to unseen samples.
Table 6: Comparison of classifiers across datasets with and without $G_{\text{btw}}$ in the centralized setting
Classifier $G_{\text{with}}$ $G_{\text{btw}}$ Dataset Train Acc Test Acc
Ridge Regression CIFAR-100 60.0 57.1
Ours CIFAR-100 60.4 57.3
Ridge Regression Imagenet-R 52.8 37.6
Ours Imagenet-R 53.4 38.6
Ridge Regression CUB200 92.0 50.4
Ours CUB200 91.3 53.7
Ridge Regression Cars 85.9 41.4
Ours Cars 86.2 44.8
Finally, we also empirically analyze in Table˜7 the impact of removing the between-class covariances in a Federated scenario using SqueezeNet. Here in the FL setup we again clearly see the negative effect of incorporating between-class scatter statistics.
Table 7: Performance comparison of different FL setups using $G_{\text{btw}}$ and $G_{\text{with}}$ across datasets.
FL Setup CIFAR-100 ImageNet-R CUB200 CARS
Using $G_{\text{btw}}$ 52.8 34.8 49.7 33.7
Using $G_{\text{btw}}$ + $G_{\text{with}}$ 56.3 36.8 51.6 42.4
Using $G_{\text{with}}$ (Ours) 56.3 37.2 53.5 44.6
Appendix G Bias of the Estimator with non-iid Client Features
In Appendix˜C we showed that, under the assumption that the per-class features are iid across clients, the proposed estimator is an unbiased estimator. In this section, we theoretically quantify the bias when the i.i.d assumption is violated.
Under the i.i.d. assumption, the single class features assigned to clients can be treated as random samples from the same population distribution with mean $\mu_{c}$ and covariance $\Sigma_{c}$. For simplicity, focusing on a single class and dropping the class subscript $c$, the population distribution has mean $\mu$ and covariance $\Sigma$. As a result, recalling Equation˜14, we can write:
To compute the expectation of the estimator $\mathbb{E}[\hat{\Sigma}$], we follow the same procedure used to prove proposition in Appendix˜C up to Equation˜13:
Substituting Equation˜21 and Equation˜19 into Equation˜20, and recalling that $N=\sum_{k=1}^{K}n_{k}$, we obtain:
where in the last step we used that $\sum_{k=1}^{K}n_{k}\mu_{k}=N\mu$.
The bias of the estimator is thus given by:
Note that if each client population covariance $\Sigma_{k}$ is equal to the global population covariance $\Sigma$, and the mean of each client $\mu_{k}$ is equal to the population mean, then the bias is zero (i.e., the estimator is unbiased). However, the bias formula reveals that when the distribution of a class within a client differs from the global distribution of the same class, our estimator introduces a systematic bias. This situation can arise in the feature-shift setting, in which each client is characterized by a different domain. We next evaluate FedCOF under the feature-shift setting to quantify how this bias affects performance in this specific scenario.
G.1 Experiments on Feature Shift Settings.
Method Acc ($\uparrow$) Comm. ($\downarrow$)
FedNCM 65.8 0.3
Fed3R 81.9 39.6
FedCOF 74.1 0.3
FedCOF (2 class means per client) 76.5 0.6
FedCOF (10 class means per client) 78.8 3.1
Following (lifedbn, ), we perform experiments with MobileNetv2 in a non-iid feature shift setting on the DomainNet (peng2019moment, ) dataset. DomainNet contains data from six different domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch. We use the top 10 most common classes of DomainNet for our experiments following the setting proposed by (lifedbn, ). We consider six clients where each client has i.i.d. data from one of the six domains. As a result, different clients have data from different feature distributions. We show in Table˜8 how training-free methods perform in feature shift settings and the accuracy to communication trade-offs.
While non-iid feature shift settings have been studied in some papers not using pre-trained models, using this setting with pre-trained models works a bit differently. When using a pre-trained model, the generalization capabilities of the pre-trained model can help in moving the distribution of class features across clients towards an iid feature distribution even if the class distribution across clients is non-iid at the image level. We believe that more comprehensive analysis of feature-shift settings when using pre-trained models requires more extensive benchmarks with higher number of clients and could be an interesting direction to explore in future works.
Appendix H Communication Costs
@@
When computing communication costs we consider that the pre-trained models are on the clients similar to (fani2024accelerating, ) and do not need to be communicated. All parameters are considered to be 32-bit floating point numbers (i.e. 4 bytes) in all our analyses and experiments.
For training-free methods, each client sends its local statistics to the server only once, as in (legate2023guiding, ; fani2024accelerating, ). In the multi-round federated learning setting, this can be efficiently implemented by ensuring on the client-side that each client only shares its local statistics during its first participation round. This avoids repeated communication of statistics from clients for all training-free methods. Following (fani2024accelerating, ), the communication from server to client is not considered since the clients does not need to receive any updates from the server as the client models are not updated in the training-free initialization of the global classifier. Thus, the communication cost is the same for a single round setting with full client participation and the setting with multiple rounds having partial client participation in each round.
Let $C_{k}$ denote the number of classes present in client $k$. Due to the non-iid Dirichlet data distribution, $C_{k}<C$, where $C$ is the total number of classes in the dataset. As a result, the communication cost varies across clients. Defining $M=\sum_{k=1}^{K}C_{k}$ as the total number of class means shared from the $K$ clients and assuming that each class mean is a $d$-dimensional feature vector, the communication cost of each evaluated approach is given by:
  • FedNCM shares a total of $M$ means from $K$ clients. Cost = $Md\cdot 32.$
  • Fed3R shares a total of $M$ sum of class features and a total of $K$ matrices of $d^{2}$ dimensionality from $K$ clients. Cost = $(Md+Kd^{2})\cdot 32.$
  • FedCOF shares a total of $M$ means from $K$ clients. Cost = $Md\cdot 32.$
  • CCVR and FedCOF-Oracle shares a total of $M$ means and $M$ covariance matrices of $d^{2}$ dimensionality from $K$ clients. Cost = $M\cdot(d+d^{2})\cdot 32.$
For the non-iid sampling in experiments reported in Table˜3, the values of $M$ for CIFAR-100, ImageNet-R, CUB200, Cars and iNat-120K are 2870, 3470, 2353, 2634 and 54590, respectively.
Instead, for federated finetuning, the communication cost is calculated as follows. Let $E$, $C$ and $H=d\cdot C$ denote the sizes of the feature extractor, the total number of classes and the classifier head, respectively. Assuming that $s<K$ is the number of clients participating in each round and $T$ is the total number of communication rounds, the communication cost for federated full-finetuning is $2\cdot(E+d\cdot C)\cdot T\cdot s\cdot 32$ and the cost for federated linear probing will be $2\cdot d\cdot C\cdot T\cdot s\cdot 32$.
For federated prompt-tuning methods (weng2024probabilistic, ), considering the classifier weights $H=dC$, $T$ rounds of communication, and $s$ clients per round, the communication costs are:
  • FedAvg-PT and FedProx-PT share the classifier weights $H=d\cdot C$ and the fixed size prompt pool $P$ for each client of size $d$ at each round. Cost = $2\cdot(d\cdot C+P\cdot d)\cdot T\cdot s\cdot 32.$
  • PFPT shares the classifier weights $H=d\cdot C$ and a variable size prompt pool $P_{i}$ for each client at each round $i$. Considering the sum of all $P_{i}$ across all $T$ rounds as $\sum_{i=1}^{T}{P}_{i}$, the communication cost is $2\cdot(d\cdot C\cdot T+d\cdot\sum_{i=1}^{T}{P_{i}})\cdot s\cdot 32.$
For the experiments reported in Table˜4, we follow the same training settings as (weng2024probabilistic, ) and train for 120 rounds with $s=10$ clients participating per round. We use a fixed prompt pool size $P=20$ for FedAvg-PT and FedProx-PT and a variable prompt pool size which is optimized over rounds for PFPT. The sums of the variable prompt pool sizes $\sum_{i=1}^{T}{P}_{i}$ across all rounds for CIFAR-100, ImageNet-R, CUB200 and Cars are 1777, 5205, 4736 and 4736 respectively.
Using PFPT (weng2024probabilistic, ) on CIFAR-100, the prompt pool size starts from 20 and ends at 10 after 120 rounds leading to improved performance and reduced communication compared to FedAvg-PT and FedProx-PT. This is consistent with the results from (weng2024probabilistic, ). However, we observe in our experiments on out-of-distribution datasets like ImageNet-R and fine-grained datasets like CUB200 and Cars that the prompt pool size increases over rounds leading to increased communication costs with respect to FedAvg-PT and FedProx-PT. We also see in Table˜4 that all existing prompt-tuning methods performs very poorly on fine-grained datasets like CUB200 and Cars.
Appendix I Improving Privacy Preservation in FedCOF
In this section we discuss additional privacy techniques to prevent the sharing of client class-wise count statistics and class means.
I.1 Adding Random Noise to Protect Class-wise Statistics.
Our method requires transmitting class-wise statistics to compute the unbiased estimator of the population covariance (Equation˜12) and classifier initialization, similar to other methods in federated learning (legate2023guiding, ; luo2021no, ). In general, transmitting the class-wise statistics may raise privacy concerns, since each client could potentially expose its class distribution. Inspired by differential privacy (10.1007/11681878_14, ), we propose perturbing the class-wise statistics of each client with different types and intensities of noise, before transmission to the global server. This analysis allows us to evaluate how robust FedCOF is to variations in class-wise statistics and whether noise perturbation mechanisms can effectively hide the true client class statistics. Specifically, we propose perturbing the class-wise statistics as follows:
$\widetilde{n}_{k,c}=\max(n_{k,c}+\sigma^{\text{noise}}_{\epsilon},0)$ (23)
We consider three types of noise:
In Figure˜6, we show that the performance of FedCOF is robust with respect to the considered noise perturbation, varying the intensity of $\epsilon\in\{0.1,0.3,0.5,0.7,0.9\}$. These results suggest that a differential privacy mechanism can be implemented to mitigate privacy concerns arising from the exposure of client class-wise frequencies. In Figure˜7, we provide a qualitative overview of how the proposed Laplacian and uniform noise perturbation affect class-wise distributions.
I.2 FedCOF with Secure Aggregation Protocols
To incorporate secure aggregation protocols with FedCOF, we propose to use pairwise masking between clients following (bonawitz2016practical, ). Each client $i$ generates a random noise vector $z_{i,j}$ for every other client $j$. Then, each of these clients $i$ computes a perturbation vector $p_{i,j}=z_{i,j}-z_{j,i}$. Similarly, $p_{j,i}=z_{j,i}-z_{i,j}$ resulting in $p_{j,i}=-p_{i,j}$. These perturbation vectors are added to the client statistics before sharing them to the server. When all client statistics are added in the server, the noise cancels out across clients and the server can use the aggregate statistics.
To use such a secure aggregation protocol with FedCOF, the communication cost will increase and become similar to Fed3R. We proceed in two steps:
  1. 1.
    Secure aggregation of class means and counts. In this phase, each client sends only class means and counts, adding perturbation vectors to the means and perturbation scalars to the counts to ensure privacy. The server then computes the global class means from the aggregated statistics and sends them back to the clients. Specifically each client $k$ sends:
    $u_{k,c}=q_{k}+n_{k,c},\qquad v_{k,c}=p_{k}+\mu_{k,c}\times n_{k,c},$
    where $n_{k,c}$ and $\mu_{k,c}$ are the original class counts and means; $p_{k}=\sum_{i\in K,i\neq k}p_{k,i}$ is the total perturbation vector for client $k$, and $\sum_{k=1}^{K}p_{k}=0$ (similarly $q_{k}$ is a perturbation scalar such that $\sum_{k=1}^{K}q_{k}=0$).
    The server aggregates these quantities, and since the perturbations cancel out, it can compute:
    $\mu_{c}=\frac{\sum_{k=1}^{K}v_{k,c}}{\sum_{k=1}^{K}u_{k,c}},\qquad N_{c}=\sum_{k=1}^{K}u_{k,c},$
    and sends $(\mu_{c},N_{c})$ back to all clients.
  2. 2.
    Secure aggregation of class-covariance terms. Each client computes
    $s_{k,c}=(N_{c}-1)n_{k,c}(\mu_{k,c}-\mu_{c})(\mu_{k,c}-\mu_{c})^{\top},$
    and forms
    $S_{k}=\sum_{c=1}^{C}s_{k,c}+M_{k},$
    where $M_{k}$ is a random matrix such that $\sum_{k=1}^{K}M_{k}=0$.
    After receiving all $S_{k}$, the server sums them, and since the noise terms cancel out, it obtains:
    $\sum_{k=1}^{K}S_{k}=\sum_{k=1}^{K}\sum_{c=1}^{C}(N_{c}-1)n_{k,c}(\mu_{k,c}-\mu_{c})(\mu_{k,c}-\mu_{c})^{\top}.$
    The final global matrix is then estimated as:
    $\hat{G}$ $=\!\frac{1}{K\!-\!1}\left(\sum_{k=1}^{K}S_{k}+\sum_{c=1}^{C}K(N_{c}-1)\gamma I_{d}\right)+N\mu_{g}\mu_{g}^{T}$
    $=\!\frac{1}{K\!-\!1}\!\left(\sum_{k=1}^{K}\sum_{c=1}^{C}(N_{c}-1)n_{k,c}(\mu_{k,c}\!-\!\mu_{c})(\mu_{k,c}\!-\!\mu_{c})^{\top}\!\!\right)\!\!+\!\frac{K}{K-1}\sum_{c=1}^{C}(N_{c}-1)\gamma I_{d}\!+\!N\mu_{g}\mu_{g}^{\top}$
    $=\!\sum_{c=1}^{C}(N_{c}-1)\left[\frac{1}{K-1}\sum_{k=1}^{K}n_{k,c}(\mu_{k,c}-\mu_{c})(\mu_{k,c}-\mu_{c})^{\top}+\gamma I_{d}\right]+N\mu_{g}\mu_{g}^{\top}.$
    where $N=\sum_{c=1}^{C}N_{c}$ and $\mu_{g}=\frac{\sum_{c=1}^{C}\mu_{c}}{N}$. This results in the exact same formulation of the matrix $\hat{G}$ as computed in the proposed FedCOF.
Although this increases training-free communication costs which becomes equivalent to Fed3R, FedCOF offers benefits in terms of accuracy and faster convergence which also reduces overall communication costs when fine-tuning after FedCOF initialization.
Appendix J Dataset and Implementation Details
Datasets. We use the following five datasets in our paper:
  • CIFAR-100 has 100 classes provided in 50k training and 10k testing images.
  • ImageNet-R (IN-R) is composed of 30k images covering 200 ImageNet classes. ImageNet-R (hendrycks2021many, ) is an out-of-distribution dataset and proposed to evaluate out-of-distribution generalization using ImageNet pre-trained weights. It contains data with multiple styles like cartoon, graffiti and origami which is not seen during pre-training.
  • CUB200 is a fine-grained dataset and has 200 classes of different bird species provided in 5994 training and 5794 testing images.
  • Stanford Cars has 196 classes of cars with 8144 training images and 8041 test images.
  • iNaturalist-Users-120k (hsu2019measuring, ) is a real-world, large-scale dataset (van2018inaturalist, ) proposed by (hsu2019measuring, ) for federated learning and contains 120k training images of natural species taken by citizen scientists around the world, belonging to 1203 classes spread across 9275 clients.
In datasets like ImageNet-R and CARS, we also face class-imbalanced situations where there is a significant class-imbalance at the global level.
@@ I.2 FedCOF with Secure Aggregation Protocols
Implementation Details. Here, we provide details on learning rate (lr) used for all fine-tuning experiments with FedAdam. For ImageNet-R and Stanford Cars, we use a lr of 0.0001 for both server and clients for FedNCM, Fed3R and FedCOF initializations. For CUB200, we use a server lr of 0.00001 and client lr of 0.00005 for Fed3R and FedCOF, while for FedNCM, we use a higher lr of 0.0001 for clients. For random classifier initialization with all datasets, we use a higher lr of 0.001 for clients and lr of 0.0001 for server. We use 1 local epoch, for all fine-tuning experiments on 4 datasets. After training-free classifier initialization, we fine-tune the models for 100 rounds. When starting from random classifier initialization, we train more for 200 rounds. When training with FedAvg and random classifier initialization, we use a client lr of 0.005 for all datasets other than inat-120K.
@@
For the linear probing (LP) experiments for the 4 datasets other than iNat-120K, with FedAvg we train for 200 rounds with 1 local epoch and use a client lr of 0.01 and server lr of 1.0 for FedNCM. For Fed3R and FedCOF initializations, we use a client lr of 0.001 and a server lr of 1.0. For LP experiments on iNat-120K, we use 3 local epochs, 30% client participation and train for 5000 rounds. For iNat-120K, we use a client lr of 0.001 for FedAvg-LP without classifier initialization, a client lr of 0.0005 for FedNCM and client lr of 0.00001 for Fed3R and FedCOF.
We use one Nvidia RTX 6000 GPU for all our experiments.
@@
The FedCOF Oracle (Sharing Full Covariances). Similar to CCVR (luo2021no, ), we aggregate the class covariances from clients as follows:
@@
For the oracle setting of FedCOF, we use the aggregated class covariance from Equation˜24 and apply shrinkage to obtain $\hat{\Sigma}_{c}+\gamma I_{d}$ and use it in Equation˜8 instead of using our estimated covariances.
Appendix K Additional Experiments
Adapting CCVR for Classifier Initialization.
Table 9: Comparison of FedCOF with CCVR across different datasets and models.
Dataset Method SqueezeNet (d=512) MobileNetv2 (d=1280) ViT-B/16 (d=768)
Acc (↑) Comm. (↓) Acc (↑) Comm. (↓) Acc (↑) Comm. (↓)
CIFAR-100 CCVR 57.5±0.2 3015.3 59.6±0.2 18823.5 72.3±0.2 6780.0
FedCOF (Ours) 56.1±0.2 5.9 63.5±0.1 14.8 73.2±0.1 8.9
IN-R CCVR 36.4±0.2 3645.7 41.9±0.2 22758.8 49.3±0.2 8197.4
FedCOF (Ours) 37.8±0.4 7.1 47.4±0.1 17.8 51.8±0.3 10.7
CUB200 CCVR 51.2±0.1 2472.1 61.6±0.2 15432.7 78.7±0.4 5558.6
FedCOF (Ours) 53.7±0.3 4.8 62.5±0.4 12.0 79.4±0.2 7.2
Cars CCVR 40.9±0.4 2767.3 36.0±0.4 17275.7 49.4±0.4 6222.5
FedCOF (Ours) 44.0±0.3 5.4 47.3±0.5 13.5 52.5±0.3 8.1
Table 10: Performance comparison of different initialization methods followed by FedAdam optimization for 100 rounds of training.
Method ImageNet-R CUB200 Cars
FedNCM+FedAdam 44.7±0.1 50.2±0.2 48.7±0.2
CCVR+FedAdam 44.6±0.3 51.5±0.2 47.9±0.1
Fed3R+FedAdam 45.9±0.3 51.2±0.3 47.4±0.4
FedCOF+FedAdam 46.0±0.4 55.7±0.4 49.6±0.6
Since CCVR (luo2021no, ) was originally proposed to calibrate classifiers after training, we adapt it to our setting and use it as an initialization method. CCVR is similar to the FedCOF Oracle in terms of communication cost as it shares class means and covariances from all clients to the server. CCVR aggregates the class covariances and means to obtain a global class distribution at the server, and then trains a linear classifier on Gaussian features sampled from aggregated class distributions from clients. We show in Table˜9 that the proposed FedCOF outperforms CCVR in most settings despite having significantly lower communication cost.
We also show in Table˜10 that FedCOF initialization is better for further fine-tuning using a pre-trained SqueezeNet that we finetune with FedAdam for 100 rounds after different initialization methods.
Linear probing after initialization experiments. We show in Figure˜8 that linear probing after FedCOF classifier initialization improves the accuracy significantly compared to FedNCM and is marginally better than Fed3R initialization across three datasets using SqueezeNet.
Refer to captionRefer to captionRefer to caption
Figure 8: Analysis of the performance with federated linear probing using FedAvg (mcmahan2017communication, ).
Table 11: Comparison of different training-free methods using SqueezeNet with training-based Fed-LP (federated linear probing with FedAvg (mcmahan2017communication, ) starting with pre-trained model and random classifier initialization) across 5 random seeds. FedNCM, Fed3R and the proposed FedCOF does not involve any training. We show the total communication cost (in MB) from all clients to server. The best results from each section are highlighted in bold.
CIFAR-100 ImageNet-R CUB200 CARS iNat-120K
@@
[…]
Comparison of training-free methods with linear probing. We also compare with our approach with the training-based federated linear probing without any initialization (where we perform FedAvg and learn only the classifier weights of models) and show in Table˜11 that FedCOF is more robust and communication-efficient compared to federated linear probing across several datasets. We follow the same settings as in Table˜3. For first 4 datasets, we perform federated linear probing for 200 rounds with 30 clients per round using FedAvg with a client learning rate of 0.01. For iNat-120k, we train more for 5000 rounds.
Impact of using pre-trained models. To quantify impact of using pre-trained models we performed experiments using a randomly initialized model and show in Table˜12 that federated training using a pre-trained model significantly outperforms a randomly initialized model using standard methods like FedAvg and FedAdam on CIFAR-10 and CIFAR-100.
Table 12: Impact of using a pre-trained SqueezeNet with different federated learning methods on CIFAR-10 and CIFAR-100. We show the total communication cost (in MB). We train a total of 100 clients with 30 clients per round for 200 rounds in non-iid settings with a Dirichlet distribution of 0.1. When starting from random initialization (no pre-training), we train for 400 rounds.
CIFAR-10 CIFAR-100
FedAvg $\times$ 37.3 74840 23.9 79248
FedAdam $\times$ 60.5 74840 44.3 79248
FedAvg $\checkmark$ 84.7 37420 56.7 39624
FedAdam $\checkmark$ 85.5 37420 62.5 39624
Experiments with ResNet-18. We perform experiments with pre-trained ResNet-18 in Table˜13. For FedAvg and FedAdam, we train for 200 rounds with 30 clients per round. For FedAvg, we train with a client learning rate of 0.001 and server learning rate of 1.0. For FedAdam, we train with a client learning rate of 0.001 and a server learning rate of 0.0001. We show that fine-tuning after FedCOF classifier initialization for 100 rounds outperforms competitive FL methods like FedAdam (which are trained for 200 rounds) by 2.5% on CIFAR-100 and 5.1% on ImageNet-R. The improved performance with FedCOF initialization validates the effectiveness of the proposed method, as it reduces communication and computation costs by half compared to FedAdam and FedAvg and still outperforms them.
Table 13: Comparison of different training-free methods using pre-trained ResNet-18 for 100 clients with training-based federated learning baselines FedAvg (mcmahan2017communication, ) and FedAdam (reddi2020adaptive, ) starting from a pre-trained model. We train for 200 rounds for FedAvg and FedAdam which uses pre-trained backbone and random classifier initialization. FedNCM, Fed3R and the proposed FedCOF do not involve any training. We also show the performance of fine-tuning with FedAdam after classifier initialization. For fine-tuning experiments we only train for 100 rounds after initialization. We show the total communication cost (in MB) from all clients to server. The best results from each section are highlighted in bold.
CIFAR-100 ImageNet-R